Q1] (20 Marks) Three-digit numbers constructed using only 0s and 1s. Let X(s) - No. of ones in the number,Y (3) - index of
Answer #1

(i)

Sample space={000,001,010,100,011,101,110,111}

P_{X,Y}(0,0)=P(\{000\})=1/8;~ P_{X,Y}(1,1)=P(\{100\})=1/8;\\P_{X,Y}(1,2)=P(\{010\})=1/8;~ P_{X,Y}(1,3)=P(\{001\})=1/8;\\ P_{X,Y}(2,1)=P(\{110,101\})=2/8;~ P_{X,Y}(2,2)=P(\{011\})=1/8;\\ P_{X,Y}(3,1)=P(\{111\})=1/8.\\\\ Joint~pmf~of~(X,Y)~is~given~as~follows:\\\\ \begin{matrix} & & & Y & & & \\ & & 0 &1 &2 & 3 &Totals & & & \\ \hline &0 & 1/8 & 0 & 0 &0 & 1/8\\ & 1 & 0& 1/8 &1/8 & 1/8 &3/8 \\ X & 2 & 0& 2/8 &1/8 &0 &3/8 \\ & 3 &0 &1/8 &0 &0 &1/8 \\\hline &Totals &1/8 &4/8 &2/8 &1/8 &1 \end{matrix}

ii.~Joint~CDF~of~X,Y~is\\\\ F_{X,Y}(x,y)=P(X\leq x,Y\leq y)\\\\ \begin{matrix} & & & Y & & & \\ & & 0 &1 &2 & 3 & \\ \hline &0 & 1/8 & 1/8 & 1/8 &1/8\\ & 1 & 1/8& 2/8 &3/8 & 4/8 \\ X & 2 & 1/8& 4/8 &6/8 &7/8 \\ & 3 &1/8 &5/8 &7/8 &1 \\\hline \end{matrix}\\\\ iii.~p_X(x)=1/8~if~x=0,3\\ =3/8~if~x=1,2\\\\ =0~otherwise\\\\ iv.~p_Y(y)=1/8~if~y=0,3\\ =4/8~if~y=1\\ =2/8~if~y=2\\ =0~otherwise